OpenStudy (anonymous):

How would you graph the formula -3y + 12x = 15

OpenStudy (cj7529):

first rearange to get y =4x - 5 and just plot three points and draw the straight line that connects them should cross the y axis at y =-5

OpenStudy (anonymous):

first you have to get in slope intercept form -3y+12x=15 -12x -12x __________________ -3y=-12x+15 divide both sides by three y =4x - 5

OpenStudy (radar):

Both sides were divided by -3 (minus three) which is good.

OpenStudy (anonymous):

 - 5 is the y intercept

OpenStudy (anonymous):

 Then you go up 4 over one to graph